# OCCAM: Online Continuous Controller Adaptation with Meta-Learned Models

### v1

@@ 6 Results
OCCAM outperforms both ablations and all baselines on all benchmarks, demonstrating the utility of both meta-learning basis functions and weight adaptation. Adaptation with our method also occurs within a few timesteps, which corresponds to 10-20 seconds of data on each system. The quadruped environment is challenging in particular, as OCCAM must predict the performance of an expressive RL policy; nonetheless, it finds behavior parameters that improve the policy’s performance. The quadrotor also presents unique challenges, as the test parameters can have up to 300% parametric error from the nominal model that must be compensated by tuning PD feedback gains (the controller does not have an integral term). These extreme parametric errors also cause instability and poor tracking in the $\mathcal{L}$1-adaptive controller. Meanwhile, our method finds gains that result in the lowest crash rate and nearly a 50% reduction in tracking error compared to the nominal gains and $\mathcal{L}$1-Adaptive controller. Both Reptile and LK-GP saturate early in the evaluations. Reptile tends to overfit to the training set and online training data [34] and gets stuck in a loop of selecting the same gains and repeatedly fitting to them.
@@ 7 Discussion and Limitations
Our results demonstrate a single algorithm that optimizes controllers on a diverse array of robotic platforms using simple search techniques. Importantly OCCAM obviates the need for manual gain tuning that typically requires domain expertise, and instead provides automatic, continual controller adaptation. This allows designers to focus on choosing performance metrics better aligned with the actual task. OCCAM is also compatible with any controller with tunable parameters, allowing straightforward application to controllers that are nonlinear, discontinuous, and not analytic. As shown in our Crazyflie experiment, OCCAM’s models and optimization routine are lightweight enough to run in realtime without a GPU.
In this work, we tested OCCAM on commonly used controllers with relatively low-dimensional gain spaces, though one might also wish to tune a higher-dimensional policy, such as a neural network. Although the policy could be conditioned on a low-dimensional behavior vector, one could also treat the weights of either the entire neural network or its last few layers [52, 53] as the controller gains and optimize those using a more powerful optimizer such as gradient descent.
Beyond this, there are a few areas in which our method could be improved. If the stochastic optimizer makes a poor initial choice and the uncertainty penalty is poorly tuned, OCCAM sometimes settles early in a suboptimal local maximum. For some systems, this incorrect choice of gains can cause crashes. Incorporating a more sophisticated optimizer and exploration strategy would be interesting future work. Secondly, if the designer has knowledge of the true system parameters at test time, there is no clear way to set the weights $w$ besides through the training data distribution. Lastly, in this work we do not rigorously evaluate how far the generalizability of our method applies, or changing the reward function $J$ online. Future work includes addressing these limitations, along with demonstrating our method on broader classes of systems.

### v2

@@ 6 Results
OCCAM outperforms both ablations and all baselines on all benchmarks, demonstrating the utility of both meta-learning basis functions and weight adaptation. Adaptation with our method also occurs within a few timesteps, which corresponds to 10-20 seconds of data on each system. The quadruped environment is challenging in particular, as OCCAM must predict the performance of an expressive RL policy; nonetheless, it finds behavior parameters that improve the policy’s performance. The quadrotor also presents unique challenges, as the parameters of the quadrotors in the test set can vary from $\frac{1}{3}\times$ to 3$\times$ the nominal parameters. In our problem setup, this must be compensated only by tuning PD feedback gains (the controller does not have an integral term). These extreme parametric errors also cause instability and poor tracking in the $\mathcal{L}$1-adaptive controller. Meanwhile, our method finds gains that result in the lowest crash rate and nearly a 50% reduction in tracking error compared to the nominal gains and $\mathcal{L}$1-Adaptive controller. Both Reptile and LK-GP saturate early in the evaluations. Reptile tends to overfit to the training set and online training data [34] and gets stuck in a loop of selecting the same gains and repeatedly fitting to them.
@@ 7 Discussion and Limitations
Our results demonstrate a single algorithm that optimizes controllers on a diverse array of robotic platforms using simple search techniques. Importantly OCCAM obviates the need for manual gain tuning that typically requires domain expertise, and instead provides automatic controller adaptation based on task-relevant metrics. OCCAM is also compatible with any controller with tunable parameters, allowing straightforward application to controllers that are nonlinear, discontinuous, and not analytic. As shown in our Crazyflie experiment, OCCAM’s models and optimization routine are lightweight enough to run in realtime without a GPU. We expect OCCAM to perform best on systems with a well-defined and relatively smooth mapping between control gains and performance; this and further assumptions are discussed in the appendix.
In this work, we tested OCCAM on commonly used controllers with relatively low-dimensional gain spaces, though one might also wish to tune a higher-dimensional policy, such as a neural network. Although the policy could be conditioned on a low-dimensional behavior vector, one could also treat the weights of either the entire neural network or its last few layers [51, 52] as the controller gains and use a more powerful optimizer.
Beyond this, there are a few areas in which our method could be improved. If the stochastic optimizer makes a poor initial choice and the uncertainty penalty is poorly tuned, OCCAM sometimes settles early in a suboptimal local maximum. For some systems, this incorrect choice of gains can cause crashes. Incorporating a more sophisticated optimizer and exploration strategy would be interesting future work. Lastly, in this work we do not rigorously evaluate how far the generalizability of our method applies, or changing the reward function $J$ online. Future work includes addressing these limitations, along with demonstrating our method on broader classes of systems.
Acknowledgments
The authors would like to thank the reviewers, who gave helpful feedback during the discussion phase. This work was supported by NSF grant CCF-2112665 (TILOS).